ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

13. Combinatorics

1991 MS Classification: 05A15, 18A99 Yoshida, Tomoyuki, Hokkaido Univ., Japan Exponential formulas and Krull-Schmidt categories For a skeletally small and locally finite category E, the generating function is defined by the formal summation

E(t) := ∑_{X} \frac{1}{|Aut(X)|} t^{X}.

Then an exponential formula

E(t) = exp (Con(E)(t))

holds if and only if the category E is a strict Krull-Schmidt category.